Questions with Complete Rationales
2026-2027 Updated.
What is Internet Savvy? - Answer -Finding reputable scientific information on the internet
-This means getting information from reliable sources, knowing where your knowledge is
coming from
Where are good places to find scientific information? - Answer NYTimes Science Section,
ScienceNews, Nautil.us (long form scientific articles), New Yorker, Atlantic Monday, Smithsonian
Magazine, National Geographic, ACADEMIC PEER-REVIEWED JOURNALS (Science, Nature, Cell
New England Journal of Medicine, The Lancel, Journal of the Medical American Association)
Google Scholar, PubMed = go to review and scroll down for external links
PubMed - Answer ▪ A service of the National Library of Medicine,National Institutes of Health
▪Searchable database of peereviewed literature
▪https://pubmed.ncbi.nlm.nih.gov/
▪Search for your topic, then you can find summary articles by clicking "review" on the left hand
side
▪Find PDFs of manuscripts by clicking the PMCID
Why is finding REPUTABLE information important? - Answer ▪Fake news is dangerous because
if you don't have the right facts you can't deal with a situation correctly
▪Not understanding how to deal with COVID because we didn't have the right information
▪ Clorox is dangerous!
▪Science builds on previous discoveries.If those discoveries are wrong, then new work will be
also in error.
Eugenics-used to justify poor treatment of enslaved people or religious minorities (current
example: Uighurs in China)
▪Climate change skepticism or denial
▪Retracted scientific paper that vaccines cause autism
Genetics Problems
Why are plants used in genetics?
,What is it called when you are looking at a single trait? - Answer -*Plants are used in genetics*
because they are *easier to breed and work with*, some even *self-fertilize*
-If you are only looking at a *single trait* (height) = *monohybrid cross*
Probability of outcome - making predictions
What happens every time you flip both coins?
The sum of all outcomes must equal? - Answer -A *coin has two sides, head and tail (like two
alleles of a gene)*. If you flip two dimes and examine the paired outcomes (whatever is face
up), you will observe three possible combinations: HH, HT, and TT.
• Since there is *half a chance that either coin will come up heads*, the probability of a
homozygous outcome, *HH (also of hh), is 1/2 x 1/2 = 1/4*
•The *heterozygous outcome*, *HT*, can arise in two ways (head/tail or tail/head) so its
*probability is 1/4 + 1/4 = 2/4*
•The outcome of many such trials of *HT x HT (flipping 2 coins and examining which pair of
faces is up)* will be *1/4HH + 2/4HT + 1/4TT*
-The *more times you flip* both coins, the *closer you will come to these ideal ratios*
-The *sum of all probable outcomes must total 1 or 100%*
A handy rule: Some of it, plus the rest of it, equals all of it.
Why use corn?
It is what kind of grass?
What is separate about them and where are they located?
What is each kernel?
An ear of corn holds what? - Answer -*Corn* (Zea mays) is a *large, coarse grass*. *Male and
female flowers are separate*. Many *male flowers* make up each of the tassels, which are
located at the *tips of stems*. The tiny pollen grains are transported by wind and make me
sneeze. *Female flowers nestle lower down*, close to the *stout part of the stem* and are
combined in structures we refer to as an ear.
•Each *kernel* in an ear of corn is the *swollen ovary of a flower*; a single strand of silk
extends from each kernel to the outside of the ensheathing husk.
-*A kernel only matures if its silk is pollinated*. Thus each kernel in an ear of corn represents a
separate fertilization of an ovum by a sperm.
-An *ear of corn* = holds the results of *hundreds of separate genetic crosses*
Genotype and phenotype
If both parents have Pp as the genotype, what would be the outcome of the monohybrid cross?
- Answer -This ear of corn illustrates the outcome of a monohybrid cross of parent plants that
were *heterozygous for kernel color* (PURPLE AND YELLOW)
•P (purple) is the dominant allele
, •p (yellow) is the recessive allele.
Both parents had the genotype Pp, and they had the purple kernel phenotype. The outcome of
Pp x Pp cross would be ¼ PP + 2/4 Pp + ¼ pp.
•There are *three different allele combination genotypes*
•There are *two diff appearance (purple or yellow) phenotypes*.
-When you predict the outcome of a cross, you want to know what you will observe, and that is
the ratio of phenotypes, purple and yellow.
The supposed monohybrid cross
What did Mendel perform and what was the varying trait?
How do you know when yellow is the recessive trait?
What are the frequencies for a phenotypic monohybrid cross? - Answer -Using the pea plant,
Gregor *Mendel originally performed monohybrid crosses* (breeding experiments that *vary
by a single trait* like pod color)
-In the original *parental generation*, one parent has *all green pods* (GG) and the other
parent has *yellow pods* (gg), upon crossing the same phenotype of green pods is expressed
for the offspring. *Green is the dominant trait of the two colors*
-When plants are allowed to self-fertilize, their progeny (F2 plants), contain some with green
pods and *others w yellow ones* which reveals that *yellow is indeed the recessive trait*
-*Both parents are presumed to be heterozygous Pp*; they had purple kernels. If this is true,
when they are crossed (Pp x Pp), you would expect the phenotypic frequencies of kernel color in
the next generation to be *¾ P_ + ¼ pp*. Is this what you observe?
•Do the ratios of each color (number of purple divided by total number of kernels counted and
number of yellow divided by total number of kernels counted) come close to your expected
phenotypic ratios?
•How can you decide if the result is close enough to support your hypothesis?
The Chi-square test
Why are phenotypes observed and expected able to differ?
What does the null hypothesis mean?
What is the chi-square test used for? - Answer -The phenotypes you observe and count
probably won't match the expected ratios exactly because *chance plays a role in biological
phenomena*, in this case, fertilization events. Chance enters again with the corn cross, since
you will not be counting all kernels on every ear, and some kernels are missing, eaten by mice.
-Is the *difference between your observation and the expected result small enough* that it
could have been produced by chance alone? This is the *null hypothesis—that there is no real
difference between the observed data and the outcome predicted by the hypothesis.*